OpenStudy (anonymous):

if we have tan(x)=2 then what is the result of (sin(x)+cos(x))/(3sin(X)-2cos(x)) using step by step solving

OpenStudy (zehanz):

tanx = 2 sinx/cosx = 2 sinx = 2cosx Substitute this in your expression: (2cosx + cosx)/(6cosx - 2cosx)= 3cosx/4cosx = 3/4 (if cosx <> 0)

OpenStudy (anonymous):

excellent answer , thanks , i like your way sloving this problem

OpenStudy (anonymous):

sinx=2cosx (2cosx+cosx)/(6cosx-2cosx) =3cosx/4cosx =3/4